OpenStudy (anonymous):

(3n^2-n)/(n^2-1)divided(n^2)/(n+1)

OpenStudy (anonymous):

factor and cancel i believe is what you need

OpenStudy (anonymous):

start with \[\frac{3n^2-n}{n^2-1}\times \frac{n+1}{n^2}\]

OpenStudy (anonymous):

then factor as \[\frac{n(3n-1)(n+1)}{(n+1)(n-1)n^2}\]

OpenStudy (anonymous):

then cancel some stuff \[\frac{\cancel{n}(3n-1)\cancel{(n+1)}}{\cancel{(n+1)}(n-1)n\cancel{n}}\]

OpenStudy (anonymous):

it was divided by, i flipped and multiplied

OpenStudy (anonymous):

just like with numbers \[\frac{4}{5}\div \frac{8}{9}=\frac{4}{5}\times \frac{9}{8}\]

OpenStudy (anonymous):

Oh ok so the answer is (3n-1)/(n-1)?